OpenStudy (anonymous):

What is 3(x+ 2/5) = x- 1/3

OpenStudy (anonymous):

x=13/30= 0.43333

OpenStudy (anonymous):

by 2/5 and 1/3 do you mean 2 divided by 5 or simply just 2/5?

OpenStudy (anonymous):

3x + 6/5=x- 1/3 2x = -1.53333 x= -.7666